OpenStudy (anonymous):

Prove that : 1/1+ x b−a+ x c−a + 1/1 + xa−b+ x c− b + 1/1+x b−c+x a −c =1 Prove that : 1/1+ x b−a+ x c−a + 1/1 + xa−b+ x c− b + 1/1+x b−c+x a −c =1 @Mathematics

jhonyy9 (jhonyy9):

1+1 +1 +xb+xb+xc+xc+xa+xa-2a-2b-2c =1+2+2xb+2xc+2xa-2a-2b-2c= =3+2x(a+b+c)-2(a+b+c) =3+(a+b+c)(2x-2)

jhonyy9 (jhonyy9):

can you prove it realy ???

OpenStudy (anonymous):

yes